Determine the margin of error for a confidence interval to estimate the population mean with nequals22 and s​ = 11.6 for the confidence levels below. ​a) 80​% ​b) 90​% ​c) 99​% ​a) The margin of error for an 80​% confidence interval is nothing.
Answer #1

Solution :

(a)

At 80%

t /2,df = 1.323

Margin of error = E = t/2,df * (s /n)

= 1.323 * (11.6 / 22)

Margin of error = E = 3.3

(b)

t /2,df = 1.721

Margin of error = E = t/2,df * (s /n)

= 1.721 * (11.6 / 22)

Margin of error = E = 4.3

(c)

t /2,df = 2.831

Margin of error = E = t/2,df * (s /n)

= 2.831 * (11.6 / 22)

Margin of error = E = 7.0